OpenStudy (dumbsearch2):

A pharmacist wants to make 30 ml of a 40% hydrogen peroxide solution. How many mls of 46% hydrogen peroxide must he mix with 36% hydrogen peroxide to make the desired solution?

hero (hero):

x + y = 30 .46(x) + .36(y) = .40(30)

hero (hero):

The first equation represents the total amount of solution after combining both mixtures. The second equation represents the total amount of hydrogen peroxide present in the solution after combining both mixtures. Solving the system yields the values of x and y. The value of x is what you will need to find for this particular problem.

OpenStudy (dumbsearch2):

Right thanks. :) I'm solving the rest, using elimination...

OpenStudy (anonymous):

what did you get?

hero (hero):

@alyygirl, Have you tried solving it?

OpenStudy (anonymous):

no

hero (hero):

Why not?

OpenStudy (anonymous):

where do i start

hero (hero):

Start by isolating y in the first equation.

OpenStudy (anonymous):

.46(x) + .36(y) = .40(30)

hero (hero):

The first equation is x + y = 30

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still stuck

hero (hero):

If x + y = 30, then y = 30 - x right?